# Balance Correlations, Agentic Zeros, and Networks: The Structure of 192 Years of War and Peace

### v1

@@ Abstract
Heider’s (1944; 1946; 1958) original P  O  X balance theory predicts human relations based on shared (or differing) perceptions and attitudes between a pair of individuals (P  O) towards an inanimate object X. Social network extensions of his theory have replaced this X with a third individual (Newcomb,, 1953; Cartwright and Harary,, 1956), leading to a plethora of adaptations that have often been inconsistent with Heider and with each other. We present a general model and formal notation for these social network extensions that permit social scientists to be more explicit about their balance theoretic statements. Specifically, we formulate statements as a comparison of two conditional probabilities of a tie, where the conditionals are defined by the 2-path relation Ego  X  Alter:
$P(Ego\stackrel{{\scriptstyle q}}{{\text{---}}}Alter\ |\ Ego\stackrel{{\scriptstyle r}}{{\text{---}}}X\stackrel{{\scriptstyle s}}{{\text{---}}}Alter)\quad>\quad P(Ego\stackrel{{\scriptstyle q}}{{\text{---}}}Alter\ |\ \neg(Ego\stackrel{{\scriptstyle r}}{{\text{---}}}X\stackrel{{\scriptstyle s}}{{\text{---}}}Alter))$
where q, r and s, represent different mutually exclusive relations. Given the importance Heider assigned to the role of negative associations, we further identify negative ties as completely separate from non-ties (neutral or zero-valued ties) and consider a signed graph to be a restricted multigraph composed of three mutually exclusive and exhaustive relations: positive ties, negative ties, and zero-ties. We stress that neutrality is the result of a triadic process. Combining these two features into our theoretical frame results in 27 identifiable configurations. Drawing on the work on Transitivity Correlation models, we propose a set of simple descriptive statistics to measure the extent to which evidence for any stipulated balance configuration is present in a network by correlating the presence of a particular tie with the corresponding conditional 2-paths that surround that tie. Finally, we demonstrate how to apply this approach to assess network-level balance in a large data set consisting of friendly vs hostile relations between countries from 1816 to 2007. We find strong evidence particularly for one of the four classic Heiderian balance theory predictions, and virtually no evidence in support of the imbalance predictions. However, we do find stable and surprising evidence that the ’neutral’ ties are important in balancing the relations among nations.
Keywords Balance Theory, Balance Correlations, Network Level Measures
@@ 1 Introduction
Balance theory prevails as one of the most widespread theories of social behavior to emerge in the past 70 years. With its roots in psychology Heider, (1958), it is widely used in sociology (e.g., Moody,, 1986; Rawlings and Friedkin,, 2017), political science (e.g., Krasner,, 1985), social psychology (e.g., Newcomb,, 1953), and anthropology (e.g., Hage,, 1979). But, perhaps no field has invested more in balance theory than social networks. Whether these networks are studied by social scientists (e.g., Doreian and Krackhardt,, 2001; Feld,, 1981), statisticians (e.g., Snijders,, 2001; Fienberg et al.,, 1985), or physicists (e.g., Kirkley et al.,, 2019), all appeal to balance theory as an explanation or motivation for actor behavior in their research.
@@ 2.1 Background
Heider (1944; 1946; 1958) originally defined balance as a function of the relationships between three entities: a person (P), another person (O), and an inanimate object (X). Each relationship between these three entities was characterized as either positive or negative. If the product of the signs is positive, the triad is deemed to be balanced; if the product is negative, the triad is deemed to be unbalanced. This yields eight combinations of possible tie patterns. The central premise in Heider’s balance theory is that P strives to attain balance due to adapting the relation to O, such that the triad becomes in any of the ’Balanced’-states shown in Table 1.
@@
- + - Balanced
- - + Balanced
+ + - Unbalanced
+ - + Unbalanced
@@
- - - Unbalanced
Table 1: Heider’s P  O  X designation.
Newcomb, (1953) was among the first to apply Heider’s theory exclusively to a triad of people. He specifically suggested that identifying X as a third individual, rather than an object as Heider had done, allowed one to use balance among three people to explain many relational social dynamics. A subsequent important paper in the development of balance theory is Cartwright and Harary (1956), which proposes formal mathematical statement about balance properties of the network as a whole, not just an isolated triple of actors. The triple was considered an unordered set without any special consideration given to one node over another — i.e., it was an unlabeled triple. However, this has obfuscated the focus on the role of “P” as the actor/agent of interest in Heider’s work. Indeed, the labels in “P  O  X” were replaced with far less discriminating labels ${i,j,k}$. This became the standard for one branch of balance theory, the concept and measure of triad census as introduced by Holland and Leinhardt, (1979) and continued to this day (e.g. Dinh et al.,, 2023).
Rapoport, (1963) translated this technical mathematical formulation into simple theoretical statements that inspired subsequent work in this area. He noted that each of these eight configurations had a social interpretation that could be construed to give these balance rules an intuition using concrete, everyday terms like friendship and enemy. He noted that the eight theoretical statements corresponding to Heider’s P  O  X model are (see Table 1):
@@
Table 2: Putting balance theory in words.
While these aforementioned simple Rapaport statements make the Heiderian predictions easy to understand, these also are inherently overly demanding. For example, no one really believes that everytime someone encounters a friend of a friend that she will establish a friendship with that person. Rather, a more reasonable prediction from this first Heiderian principle is that a friend of a friend is more likely to be a friend than if that person is not a friend of a friend. That is, the 2-path relation, friend-of-a-friend, creates a favorable context in which a friendship is more likely to form.
@@ 2.2 Balance theory as a set of probability statements
We will throughout this paper replace Heider’s P  O  X terms with corresponding network terms. We start by noting that the only relationship that Heider specifically designated as interpersonal was the relationship between P and O. And the question posed by the model is what relation P would choose to have with O, given the association both have with X. This X provides a contextual element that influences P’s choice. Consistent with the social network literature, we refer to this P  O relationship as an Ego  Alter relationship, wherein Ego is choosing a relationship with Alter. We retain the label “X” to refer to a contextual third party. With this minor amendment, we can formalize the first Heiderian “a friend of a friend is a friend” principle with the following conditional inequality:
$\hskip 36.135ptP(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ Ego\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter)>P(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ \neg(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter))\hss$ (1)
or, equivalently:
$\hskip 36.135ptP(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ Ego\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter)-P(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ \neg(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter))>0\hss$ (2)
where $\stackrel{{\scriptstyle f}}{{\text{---}}}$ indicates a friendship tie. Similar conditional inequalities can be made of each of the other three primary Heiderian predictions about balanced relations. In cases of imbalanced relations, such as “a friend of an enemy is a friend,” then the prediction is that the 2-path relation, being a friend of an enemy, should lead to relatively infrequent friendships. Such imbalanced statements would be similarly described as in (1), except that the inequality is reversed:
$\hskip 36.135ptP(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ Ego\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter)<P(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ \neg(Ego\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter))\hss$ (3)
or, equivalently:
$\hskip 36.135ptP(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ Ego\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter)-P(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ \neg(Ego\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter))<0\ \hss$ (4)
@@
At the network level, we can frame this prediction as a statement about the relative likelihood that the predicted tie would appear given that it is facing a contextual 2-path relation. That is, in the network as a whole, we can cross-tabulate the number of occurrences of a tie type with the number of occurrences of the specified 2-path relation:
@@
Table 3: Cross-tabulation of Ego  Alter ties with 2-path triples, Ego  X  Alter. A, B, C and D are counts of co-occurrences.
@@
If we assume a uniform and random draw from all $\{i,j,k\}$ triples in the network, then we can directly calculate the relevant conditional probabilities from this 2x2 table. The first conditional in Eq. (1) and (2), the probability that a friendship between Ego and Alter exists, given that there is an X who is also a friend of Ego and a friend of Alter, is calculated as $\frac{A}{A+C}$. The second conditional in each equation are calculated as $\frac{B}{B+D}$. Thus, the difference between these two fractions $\frac{A}{A+C}-\frac{B}{B+D}$ represents the extent to which the particular balance theory prediction is true. The larger the difference, the stronger the effect size given the conditional 2-path. If the the relation is unbalanced, such as described in Eq. 3 and 4, then the difference between these two fractions would be negative. The larger the negative value, the stronger the effect of the 2-path in suppressing the existence of the Ego  Alter tie.
@@
Conversely, if $\frac{A}{A+C}-\frac{B}{B+D}$ is positive, then it will also be the case that the difference between the corresponding row percentages will be positive (i.e., $\frac{A}{A+B}-\frac{C}{C+D}>0$). This represents the symmetric claim that if the Ego  Alter tie exists, then the probability of the designated 2-path Ego  X  Alter tie will be greater than if the Ego  Alter tie does not exist. While these differences will not generally equal each other, they will always be of the same sign. Moreover, the size of these differences will correspond to the effect size of the particular Heiderian balance principle.
@@
Putting this altogether, we can describe the direction and size of the balance theory effect by taking the geometric mean of these two differences in probabilities, these two descriptors of balance in the network:
$\sqrt{\left(\frac{A}{A+C}-\frac{B}{B+D}\right)\left(\frac{A}{A+B}-\frac{C}{C+D}\right)}$
@@
Finally, we draw on earlier work Dekker et al., (2019) to note that this geometric mean is equal to the Pearson correlation between the vector of their corresponding Ego  X  Alter 2-path ties for all $\{i,j,k\}$ triples and the corresponding vector of their associated Ego  Alter ties. That is,
$\rho=s\sqrt{\left(\frac{A}{A+C}-\frac{B}{B+D}\right)\left(\frac{A}{A+B}-\frac{C}{C+D}\right)}$ (5)
where:
$s=sgn\left(\frac{A}{A+C}-\frac{B}{B+D}\right).$ (6)
@@
The sign function in 6 differentiates between balanced and unbalanced predictions. If $\frac{A}{A+C}>\frac{B}{B+D}$, then the correlation between the presence of the predicted Ego–Alter tie and its conditional 2-path will be positive; thus, the positive root of this square root function is used, indicating that the Ego  Alter tie probability is enhanced in the presence of the Ego  X  Alter 2-path. If $\frac{A}{A+C}<\frac{B}{B+D}$, then this correlation will be negative; thus, the negative root of this square root function is used, indicating that the Ego  Alter tie probability is suppressed in the presence of the Ego  X  Alter 2-path.
@@
In other words, the Pearson product-moment correlation between observed Ego  Alter ties and their associated Ego  X  Alter 2-path ties directly measures the extent to which balance (or unbalance) is observed in the network.
@@ 2.3 The Critical Role of (Agentic) Non-ties
Another restriction in Heider’s model is that it assumes all relations are either positive or negative, not neutral. On the other hand, as Cartwright and Harary noted (1956) very early on, in reality many people are going to have neutral sentiments (neither positive nor negative) about each other. While such non-ties are often seen as part of any natural network, they are often treated as “missing data” and are excluded from any predictions (Cartwright & Harary, 1956). Or, by some conventions, as noted by Rawlings and Friedkin (2017, p. 515), “… the absence of a positive sentiment implies the presence of a negative sentiment,” and as such a lack of a tie may be deemed to represent a negative tie. We disagree completely.
In a major departure from this tradition, we explicitly consider neutral relations as agentic or social choices, just as much as positive and negative relations. There are many reasons that neutral or non-relationship ties should be considered within a balance theory context. For example, in a highly polarized and conflict-ridden environment, Ego may have positive sentiments toward X, but Ego also observes that Alter and X have an intense dislike for each other. One way of escaping this uncomfortable position is to establish a neutral or non-relationship with Alter, rather than taking sides in the dispute. That is, Ego may take the “Switzerland” option of staying neutral in a spat, in order to preserve his options to work with Alter in the future.
Strategic neutrality is not the only reason for considering neutral relations in balance theory. Within the blockmodeling perspective it has been noted frequently how zero blocks (groups of people who are discouraged from relating to each other or to other groups) can be found as a way to preserve social order (White, Boorman & Breiger, 1976; Pachucki & Breiger, 2010). The cost of forming ties under such conditions could be substantial in a social situation. Understanding how these non-ties play a role in balancing triadic relations is explicitly part of our model.

### v2

@@ Abstract
Social network extensions of Heider’s classic balance theory have led to a plethora of adaptations that have often been inconsistent with Heider and with each other. Structural balance theory as one of the branches of balance theory has mainly focused on graph partitioning, thereby assuming, for example, homogeneity in balance-driven behavior of nodes over time. We present a general model and formal notation for these social network extensions that permit social scientists to be more explicit about their balance theoretic statements, including testing of the behavioral assumptions. Specifically, we formulate statements as a comparison of two conditional probabilities of a tie, $Ego\stackrel{{\scriptstyle q}}{{\text{---}}}Alter$ , where the conditionals are defined by the 2-path relations $Ego\,\stackrel{{\scriptstyle r}}{{\text{---}}}\,X\,\stackrel{{\scriptstyle s}}{{\text{---}}}\,Alter$ and the negation, $\neg(Ego\,\stackrel{{\scriptstyle r}}{{\text{---}}}\,X\,\stackrel{{\scriptstyle s}}{{\text{---}}}\,Alter)$. The key here is that q, r and s, are indices that identify elements in a set of mutually exclusive and exhaustive relations. In other words, this approach releases researchers of the confining assumption of a signed graph dichotomy. Here we identify neutral or non-ties as distinct from negative ties by converting a signed graph to be a restricted multigraph composed of three mutually exclusive and exhaustive relations: positive ties, negative ties, and neutral-ties. Drawing on the work on Transitivity Correlation models, we develop a set of straightforward descriptive statistics to measure the prevalence for any stipulated balance configuration in a network by correlating the presence of a particular tie with the corresponding conditional 2-paths that surround that tie. Two major advantages are that such balance correlations can be compared directly even if network sizes and densities differ, and that specific (un)balance behaviors can be evaluated. We apply this approach to assess network-level balance in a large data set consisting of friendly vs hostile relations between countries from 1816 to 2007. We find strong evidence for one of the four classic Heiderian balance theory predictions, and virtually no evidence in support of the unbalanced predictions. However, we do find stable and surprising evidence that the “neutral” ties are important in balancing the relations among nations. Results further suggest that prevalence of balance driven behavior varies over time, and in fact other triadic motivated behaviors prevail among countries in certain eras.
@@ 1 Introduction
Balance theory prevails as one of the most widespread theories of social behavior to emerge in the past 70 years. With its roots in psychology [26], it is widely used in sociology [34, 42], political science [32], social psychology [36], and anthropology [21]. But, perhaps no field has invested more in balance theory than social networks. Whether these networks are studied by social scientists [12, 16], statisticians [45, 18], or physicists [31], all appeal to balance theory as an explanation or motivation for actor behavior in their research.
@@
Most notably we want to address the gap that has risen since [6] proved a result on network partitioning as a consequence of balance seeking behavior of nodes. This shifted research away from certain crucial elements in balance theory as it was originally conceived. Already, [37] noticed different strands of balance theory had come about. Where social network theory mostly focuses on structural balance theory as an explanation for group formation and partitioning in networks, the social psychology strand considers diversity in nodal affect and behavior underlying balance processes [37, 11, 14]. What is not considered in neither branch is the prevalence of such behavior in the whole signed network. This is fundamental to understanding the micro-macro link, and this paper explores this issue for the network of international relations over nearly 200 years. Measuring and testing behavioral assumptions of balance theory requires to release the dichotomous relation restriction, and hence requires another more general model.
@@ 2.1 Background
Heider [24, 25, 26] originally defined balance as a function of the relationships between three entities: a person (P), another person (O), and an inanimate object (X). Each relationship between these three entities was characterized as either positive or negative. If the product of the signs is positive, the triad is deemed to be balanced; if the product is negative, the triad is deemed to be unbalanced. This yields eight combinations of possible tie patterns. The central premise in Heider’s balance theory is that P strives to attain balance due to adapting the relation to O, such that the triad becomes in any of the ’Balanced’-states shown in Table 1.
Table 1: Heider’s P  O  X designation.
@@
-- -- $+$ Balanced
$+$ $+$ -- Unbalanced
$+$ -- $+$ Unbalanced
@@
-- -- -- Unbalanced
Newcomb [36] was among the first to apply Heider’s theory exclusively to a triad of people. By identifying “X" as a third individual, rather than an object as Heider had done, he demonstrated how we could use balance among three people to explain many relational social dynamics. An important extension of this idea was developed by Cartwright and Harary [6], who proposed formal mathematical statements about balance properties of the network as a whole, not just isolated triples of actors. The triple was considered an unordered set without any special consideration given to one node over another — i.e., it was an unlabeled triple. However, this has obfuscated the focus on the role of “P” as the actor/agent of interest in Heider’s work. Indeed, the labels in “P  O  X” were replaced with far less discriminating labels ${i,j,k}$. This became the standard for one branch of balance theory [37, 11, 38] (see those for description of different branches), the concept and measure of triad census as introduced by [27] and continues to this day [10, 2, 3, 15, 14].
Rapaport [41] provided an important theoretical contribution by translating these technical and mathematical elements into simple statements of social relations that has inspired much of the subsequent work in this area. Specifically, each of these eight configurations has a social interpretation that could be construed to give these balance rules an intuition in concrete, everyday terms like friendship and enemy. These eight theoretical statements corresponding to Heider’s P  O  X model are in Table 2.
Table 2: Putting balance theory in words.
@@
While such statements make the Heiderian predictions easy to understand, these also are inherently overly demanding. For example, no one really believes that everytime someone encounters a friend of a friend that she will establish a friendship with that person. Rather, a more reasonable prediction from this first Heiderian principle is that a friend of a friend is more likely to be a friend than if that person is not a friend of a friend. That is, the 2-path relation, friend-of-a-friend, creates a favorable context in which a friendship is more likely to form than without that context. Unfortunately, measures of the Degree of Balance, usually don’t allow to distinguish between these different types of (un)balance behaviors [6, 22, 9, 2, 15].
@@ 2.2 Balance theory as a set of probability statements
We will throughout this paper replace Heider’s P  O  X terms with corresponding network terms. We start by noting that the only relationship that Heider specifically designated as interpersonal was the relationship between P and O. And the question posed by the model is what relation P would choose to have with O, given the association both have with X. This X provides a contextual element that influences P’s choice. Consistent with the social network literature, we refer to this P  O relationship as an Ego   Alter relationship, wherein Ego ($E$) is choosing a relationship with Alter ($A$). We retain the label “X” to refer to a contextual third party. With this minor amendment, we can formalize the first Heiderian “a friend of a friend is a friend” principle with the following conditional inequality:
$P(E\stackrel{{\scriptstyle f}}{{\text{---}}}A\ |\ E\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}A)>P(E\stackrel{{\scriptstyle f}}{{\text{---}}}A\ |\ \neg(E\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}A))$ (1)
or, equivalently,
$P(E\stackrel{{\scriptstyle f}}{{\text{---}}}A\ |\ E\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}A)-P(E\stackrel{{\scriptstyle f}}{{\text{---}}}A\ |\ \neg(E\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}A))>0$ (2)
where $\stackrel{{\scriptstyle f}}{{\text{---}}}$ indicates a friendship tie. Similar conditional inequalities can be made of each of the other three primary Heiderian predictions about balanced relations. In cases of unbalanced relations, such as “a friend of an enemy is a friend,” then the prediction is that the 2-path relation, being a friend of an enemy, should lead to relatively infrequent friendships. Such unbalanced statements would be similarly described as in 1, except that the inequality is reversed:
$P(E\stackrel{{\scriptstyle f}}{{\text{---}}}A\ |\ E\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}A)<P(E\stackrel{{\scriptstyle f}}{{\text{---}}}A\ |\ \neg(E\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}A))$ (3)
or, equivalently,
$P(E\stackrel{{\scriptstyle f}}{{\text{---}}}A\ |\ E\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}A)-P(E\stackrel{{\scriptstyle f}}{{\text{---}}}A\ |\ \neg(E\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}A))<0$ (4)
@@
At the network level, we can frame this prediction as a statement about the relative likelihood that the predicted tie would appear given that it is facing a contextual 2-path relation. That is, in the network as a whole, we can cross-tabulate the number of occurrences of a tie type with the number of occurrences of the specified 2-path relation, shown in Table 3.
Table 3: Cross-tabulation of Ego  Alter ties with 2-path triples, Ego  X  Alter.
@@
Note: Co-occurrence type counts are given by $a$, $b$, $c$, and $d$.
@@
If we assume a uniform and random draw from all $\{i,j,k\}$ triples in the network, then we can directly calculate the relevant conditional probabilities from this 2x2 table. The first conditional in Eq 1 and Eq 2, the probability that a friendship between Ego and Alter exists, given that there is an X who is also a friend of Ego and a friend of Alter, is calculated as ${a}/{(a+c)}$. The second conditional in each equation is calculated as ${b}/{(b+d)}$. Thus, the difference between these two fractions $({a}/{(a+c)})-({b}/{(b+d)})$ represents the extent to which the particular balance theory prediction is true. The larger the difference, the stronger the association given the conditional 2-path. If the relation is unbalanced, such as described in Eq. 3 and 4, then the difference between these two fractions would be negative. The larger the negative value, the stronger the effect of the 2-path in suppressing the existence of the Ego  Alter tie.
@@
Conversely, if $({a}/{(a+c)})-({b}/{(b+d)})$ is positive, then it will also be the case that the difference between the corresponding row percentages will be positive (i.e., $({a}/{(a+b)})-({c}/{(c+d)})>0$). This represents the symmetric claim that if the Ego  Alter tie exists, then the probability of the designated 2-path Ego  X  Alter tie will be greater than if the Ego  Alter tie does not exist. While these differences will not in general exactly equal each other, they will always be of the same sign. Moreover, the size of these differences will correspond to the strength of the particular Heiderian balance principle.
@@
Putting all this together, we can describe the size of the balance theory effect by taking the geometric mean of these two differences in probabilities:
$\sqrt{\left(\frac{a}{a+c}-\frac{b}{b+d}\right)\left(\frac{a}{a+b}-\frac{c}{c+d}\right)}.$
@@
Finally, we draw on earlier work [8] to note that this geometric mean is equal to the Pearson correlation between the vector of their corresponding Ego  X  Alter 2-path ties for all $\{i,j,k\}$ triples and the corresponding vector of their associated Ego  Alter ties. That is,
$\rho=s\sqrt{\left(\frac{a}{a+c}-\frac{b}{b+d}\right)\left(\frac{a}{a+b}-\frac{c}{c+d}\right)},$ (5)
where
$s=\operatorname{sign}\left(\frac{a}{a+c}-\frac{b}{b+d}\right).$ (6)
@@
The sign function in Eq 6 differentiates between balanced and unbalanced predictions. If ${a}/{(a+c)}>{b}/{(b+d)}$, then the correlation between the presence of the predicted Ego–Alter tie and its conditional 2-path will be positive; thus, the positive root of this square root function is used, indicating that the Ego  Alter tie probability is enhanced in the presence of the Ego  X  Alter 2-path. If ${a}/{(a+c)}<{b}/{(b+d)}$, then this correlation will be negative; thus, the negative root of this square root function is used, indicating that the Ego  Alter tie probability is suppressed in the presence of the Ego  X  Alter 2-path.
@@
In other words, the Pearson product-moment correlation between observed Ego  Alter ties and their associated Ego  X  Alter 2-path ties directly measures the observed prevalence of balance (or unbalance) in the network as a whole.
@@ 2.3 The critical role of (agentic) non-ties
Another restriction in Heider’s model is that it assumes all relations are either positive or negative, not neutral. On the other hand, in reality many people are going to have neutral sentiments (neither positive nor negative) about each other. While such non-ties are often seen as part of any natural network, they are often treated as “missing data” and are excluded from any predictions [6]. Or, by some conventions, as noted by [42], “… the absence of a positive sentiment implies the presence of a negative sentiment,” and as such a lack of a tie may be deemed to represent a negative tie.
In a major departure from this tradition, we explicitly consider neutral relations as agentic or social choices, just as much as positive and negative relations. There are many reasons that neutral or non-relationship ties should be considered a choice within a balance theory context. For example, in a highly polarized and conflict-ridden environment, Ego may have positive sentiments toward X, but Ego also observes that Alter and X have an intense dislike for each other. If Ego were to develop a positive relation with Alter, even if it were to serve Ego’s interests to do so, this would create an unbalanced triple. One way of escaping this uncomfortable position is to establish a neutral or non-relationship with Alter, rather than taking sides in the dispute. That is, Ego may take the “Switzerland” option of staying neutral in the spat, in order to preserve her options to work with Alter in the future.
Strategic neutrality is not the only reason for considering neutral relations in balance theory. Within the blockmodeling perspective it has been noted frequently how zero blocks (groups of people who are discouraged from relating to each other or to other groups) can be found as a way to preserve social order [46, 39]. The cost of forming ties under such conditions could be substantial in a social situation. Understanding how these non-ties play a role in balancing triadic relations is explicitly part of our model.
Yet another reason for observing null ties, even in cases where relationships to all other actors are evaluated is a limitation in resources. Such restrictions prevent engagement in positive or negative ties. For example, in international relations smaller countries may not have sufficient budget or human capital to invest in embassies in each other country. Also, an isolationist approach to foreign affairs may prevent joining collective security agreements. Or negatively, countries cannot afford to go at war in case of resource dependencies. Although different ideologies, internal policies and choices of countries may lead to neutrality, the behavioral consequence is uniform, and distinct from positive and negative ties.